NASA ISS On-Orbit Status 22 April 2014

©NASA

NASA ISS On-Orbit Status 22 April 2014.

Today: Extravehicular Activity (EVA) #26 Preparation: In preparation for tomorrow's planned EVA to Remove and Replace (R&R) the failed External (EXT) Multiplexer/Demultiplexer (MDM)2, the crew performed the following tasks:

Flight Engineer (FE)-3 Swanson and FE-5 Mastracchio performed final tool configuration.
Commander (CDR) Wakata completed a tool audit.
Wakata, Swanson and Mastracchio completed Dynamic Onboard Ubiquitous Graphics (DOUG) software review.
All 3 crew members reviewed EVA procedures.
Wakata completed part 2 of Airlock preparation to support EVA operations.
Hatch opening is scheduled for 8:25am CDT.

NanoRacks BioRacks: Following installation yesterday the ground was unable to receive data from the NanoRacks BioRack. Troubleshooting was unsuccessful and today's Experiment Container and Controller installation was delayed and Nanoracks Platform-3 was powered off. Swanson performed additional troubleshooting steps this morning with no success. Ground teams are working on a forward plan to recover the BioRack hardware.

T-Cell Finalization: Mastracchio copied and transferred the T-Cell data from the KUBIK-3 to the European Drawer Rack (EDR) laptop. He then disconnected and stowed the KUBIK-3. T-Cell samples will be returned to the ground for analysis on Dragon at the end of the SpaceX-3 mission. T-Cell seeks the cause of a depression in the human immune system while in microgravity. T-cells, a type of white cell, are coated with chemical receptors that must trigger together to activate the body's immune system properly. T-cells from flight crews and ground volunteers in a range of ages are analyzed to determine changes in protein production and gene response on the ground and in microgravity.

SpaceX (SpX)-3 Status: Dragon systems are nominal. Powered cargo still on board is nominal (General Laboratory Active Cryogenic ISS Experiment Refrigerator [GLACIER], Optical PAyload for Lasercomm Science [OPALS], and High Definition Earth Viewing [HDEV]). Dragon cargo operations are approximately 20% complete.

Today's Planned Tasks

SM ПСС (Caution & Warning Panel) Test [Complete]
Laptop RS1(2) Reboot / Onboard Computer System (БВС), [Complete]
RSS1,2 Reboot / r/g 1009 [Complete]
Daily Planning Conference (S-band) [Complete]
Pre-EVA Crew Health Status - Prep [Complete]
Final printout of EVA procedures [Complete]
GLACIER - Hardware Deployment [Complete]
Video Footage of Greetings / r/g 5444 [Complete]
Nanoracks/Biorack troubleshooting [Complete]
Nanoracks/Biorack install [Aborted]
Pre-EVA Crew Health Status - CMO [Complete]
Pre-EVA Crew Health Status - Subject [Complete]
Dismantle ЛКТ (ТА251М1Б) and ПЗУ in ТГК 421 [Complete
HDPCG. Activation and MELFI Insertion [Complete]
EVA Tool Config [Complete]
Study of veins in lower extremities r/g 5437 [Complete]
On MCC GO Urine Transfer from EDV-U to Progress 421 (AO) Water Tank 1 / r/g 5442 [Complete]
MRM2 comm config to support voice comm during the experiment [Complete]
Kulonovskiy Kristall Experiment Run r/g 5440 [Complete]
EVA Procedure Conference [Complete]
Pre-EVA Crew Health Status - CMO [Complete]
Pre-EVA Crew Health Status - Subject [Complete]
Pre-EVA Crew Health Status Data Entry and Stow [Complete]
IMS Update [Complete]
DOUG Software Review [Complete]
Restore Nominal Comm Configuration [Complete]
ТКГ 427 (DC1) Transfers and IMS Ops / ТКГ 427 Cargo Transfers (РПР) + r/g 5317 [Complete]
KULONOVSKIY KRISTALL. Copy and Downlink Data via РСПИ / r/g 5440 [Complete]
EVA2 Procedure Review [Complete]
Gas Analyzer О2 channel adjustment using Analyzer Status Indicator [Complete]
ТКГ 421 (AO) Activation, Air Duct Removal / Deactivation/Activation [Complete]
СОЖ Maintenance [Complete]
EVA Procedure Conference [Complete]
PELLE Download & placement before EVA / r/g 5438 [Complete]
Connecting HDV Sony HVR-Z7E camcorder and Activation of MPEG2 Multicast TV Signal Monitoring. [Complete]
Quick Release Screw Clamps Removal from AO side, Video of the interface [Complete]
DCB - Coldbag Unpack [Complete]
MICRO7 - Fixation Kit Check [Complete]
Cleaning camera digital image sensors. r/g 5439 [Complete]
Crew Preference Kit (КСПЭ) hardware Deactivation and closing applications [Complete]
Closure of SM Aft - ТКГ 421 Transfer Hatches / Deactivation/Activation [Complete]
On MCC GO ПрК-СУ and ТКГ-СУ ТГК 421 (АО) Hatch Leak Check [Complete]
Equipment Lock Preparation [Complete]
Disconnect Integrated Command and Telemetry System (ЕКТС) АФУ cable from PCE АФУ r/g 5441 [Complete]
TCELL. Copying Temperature and Centrifuge Data [Complete]
NANO - Module-9 Tube Insertion in CGBA [Complete]
EVA Tool Config [Complete]
TCELL. Copy KUBIK 3 data files [Complete]
Audit of space behind ISS RS panels / r/g 5443 [Complete]
EDR KUBIK3 deactivation and stowage after the experiment [Complete]
Task List

ARED Cylinder Evac
IFM Sabatier purge [Complete]
WHC KTO Replace
WHC EDV R&R
Roscosmos Report
Uragan Photo
Ekon Photo
HMS-OHA test (CDR) [Complete]
HMS-OHA test (FE-5)
WHC UR/IF R&R
No skid tape audit
ARED SSCV5 pre-gather
Data prep return
CSA-CP C/O P2
KTO EDV audit
MI core return prep [Complete]
ST-ISS read
XFER Dragon Cargo Ops
Ground

Nominal commanding
Three-Day Look Ahead:

Wednesday, 04/23: 53P undock, EVA 26
Thursday, 04/24: Post EVA 26 cleanup, crew day off
Friday, 04/25: 53P redock, MICRO-7, Sprint Ultrasound, Hybrid Training, Skin B
QUICK ISS Status - Environmental Control Group:

Component - Status
Elektron - Off
Vozdukh - Manual
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Stop
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Standby
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up

Please follow SpaceRef on Twitter and Like us on Facebook.